A bus travels 125 miles in the same time an airplane travels 1515 miles. If the rate of the car is 315 miles per hour less than the rate of the airplane, what is the rate of each? Let r represent the rate of the car.
Answer by lwsshak3(11628)

A bus travels 125 miles in the same time an airplane travels 1515 miles. If the rate of the car is 315 miles per hour less than the rate of the airplane, what is the rate of each? Let r represent the rate of the car.
**
let r=rate of car
r+315=rate of airplane
Travel time = distance/rate of speed (same for both)
125/r=1515/r+315
cross-multiply
125r+125*315=1515r
1515r-125r=125*315
1390r=39375
r=28.32
r+315=343.33
ans:
rate of speed of car=28.32 mph
rate of speed of airplane=343.33 mph